OpenStudy (anonymous):

Solve: −3(x − 2) = 15 My answer is -7 Am I right?

OpenStudy (anonymous):

-3x+6 = 15 -3x = 9 x = -3

OpenStudy (anonymous):

answer is -3

OpenStudy (anonymous):

Oh ok, Thank you! :)

OpenStudy (anonymous):

Oh I see it now! Thanks so much!